OpenStudy (anonymous):

Find the distance (in miles) along an arc on the surface of the earth that subtends a central angle of 3 minutes (1 minute = 1/60 degree). The radius of the earth is 3960 miles.

OpenStudy (dan815):

s0 3/60 degrees is what we want arc of

OpenStudy (dan815):

the radius is 3960 miles so the whol circumference is Circum=2pi*r=2*pi*3960 now for 360 degrres = full circumference 60 degrees is 1/6th of 360=1/6th of full cirfumference 3/60 = 3/60^2 of 60=3/60^2*1/6h of full circumference

OpenStudy (dan815):

or just do this

OpenStudy (dan815):

360 deg = 2pi*r 3/60 deg = x ---------- 120*60*3/60=120*60*x 360=7200x 360 =2*pi*r so 7200 x= 2pi*r x=2*pi*r/7200

OpenStudy (dan815):

im seeing what factor to multiply the 2nd equation to make it 360 deg = something

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

the answer is 3.455751919 miles